OpenStudy (anonymous):

Identify the equation and graph of the line passing through point (2, -4) and having a slope of 3/2

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

3x-2y=14 3x+2y=-2 -2x+3y=-16 2x+3y=-8

OpenStudy (e.mccormick):

Askin Mathematics, not computers.